# From chimeras to extensive chaos in networks of heterogeneous Kuramoto oscillator populations

### v1

@@ IV Review of the two-population model
Figure 2: Bifurcation diagrams of the two population model (Eq. (3) with Eq.(16)) for $\alpha=1.2$ and $\Delta=1$. (a) Two-parameter bifurcation diagram. Solid lines indicate supercritical bifurcations (+) and dashed lines indicate subcritical bifurcations (-). Black curves correspond to the pitchfork bifurcation (PF) of the homogeneous state (given by Eq. (14)), red curves indicate Hopf bifurcations (H), green curves indicate double homoclinic (DHom.) and heteroclinic (Het.) bifurcations, the blue curve indicates a saddle-node bifurcation (SN), and the purple curve indicates a saddle-node of limit-cycles (SNLC). Straight dotted thin line indicates the Kuramoto synchronization transition (H0). Color-shaded regions indicate instability of the homogeneous solution, with each color corresponding to a different attractor set: red for stable steady chimeras, yellow for breathing chimeras, green for symmetric oscillations, blue for bistability between symmetric oscillations and antiphase state, and purple for stability of the antiphase fixed point only. (b,c) One-parameter bifurcation diagram for $\mu=0.9$. Solid curves with different colors indicate different attractor types: Black for stable fixed points, grey for unstable fixed points, red for stable limit-cycles, and light red for unstable limit-cycles. Relevant bifurcations are labelled on the top of the figures, with the different regions between bifurcations shaded with the same color scheme as in panel (a). Most of these curves have been obtained with auto-07pDoedel et al. (2007) (see Appendix B).
As a first example we revisit the two-population model. Previous works on this setup showed the emergence of chimera and other states from a pitchfork bifurcation of the homogeneous synchronized state if $\Delta>0$Laing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021). Here we show that such bifurcation corresponds to a transverse instability provided by Eq. (14), and review some of the dynamics emerging from this symmetry-breaking. In order to simplify the analysis, we fix $\Delta=1$ without loss of generality (see Appendix C). This choice is purposely different from that used in previous studies in order to highlight that "weak heterogeneity" is not a requirement for the emergence of this rich bifurcation scenario. Similarly, we fix $\alpha=1.2$, a value relatively far from the synchronization transition at $\alpha=\pi/2$.
@@
The eigenvalues of this connectivity matrix read $\Lambda_{1}=1$ and $\Lambda_{2}=2\mu-1$. Substituting $\Lambda_{2}$ in Eq. (14) and setting $\text{Re}[\lambda_{2}^{(+)}]=0$ provides an analytical expression for the pitchfork bifurcation of the homogeneous state as an implicit equation on the system parameters. Black curve in Figure 2(a) shows such bifurcation line in the $(K,\mu)$-plane, with the color-shaded region indicating the region of instability (see also black curve in 1(b)). Within this region, four different states might exist (apart from the incoherent and coherent homogeneous states) [^5] For $\alpha=1.2$ we did not find the chaotic state unveiled in [Laing, 2012], but it can be attained, for instance, at $(K,\mu,\alpha,\Delta)=(20,0.9,1.25,1)$:
  • Steady chimera, a fixed point characterized by $R_{1}\neq R_{2}$.
  • Breathing chimera, a limit-cycle characterized by one population having always a larger degree of synchrony than the other, i.e., either $R_{1}(t)>R_{2}(t)$ or $R_{2}(t)>R_{1}(t)$ for all $t>0$ (see Fig. 3(a)).
  • Symmetric limit-cycle, characterized by $R_{1}(t)=R_{2}(t+T/2)$, where $T$ is the period of the oscillation (see Fig. 3(b)).
  • Antiphase fixed point, characterized by $R_{1}(t)=R_{2}(t)$ but $\Phi_{1}-\Phi_{2}=\pi$. A thorough analysis of this state, including linear stability, is provided in Appendix D.
In order to illustrate the emergence and nature of these different states, we first fix $\mu=0.9$ and study the system dynamics upon increasing $K$ with the help of numerical continuation softwareDoedel et al. (2007). The resulting bifurcation diagram is depicted in Figs 2(b) and (c) and can be split in nine different regions (labelled i-ix) delimited by different bifurcations:
@@
    For $K<K_{c}\approx 5.52$ the incoherent state $Z_{\sigma}=Z^{(0)}=0$ is the only stable solution. This state loses stability at the Kuramoto synchronization transition $H_{0}$, leading to the emergence of the coherent homogeneous state.
@@
    The coherent homogeneous state is the global attractor of the system from $H_{0}$ up to $K\approx 6.66$, when it loses stability through the supercritical pitchfork bifurcation (PF+) given by Eq. (14).
@@
    From PF+, two stable fixed points emerge, corresponding to steady chimeras (black lines in Fig. 2(b), red-shaded region). Although the steady chimeras coexist with the unstable homogeneous and antiphase fixed points, they are the sole attractor in this region of the parameter space.
@@
    At $K\approx 7.37$ the steady chimeras undergo a supercritical Hopf bifurcation (H+) leading to the emergence of two stable limit-cycles, corresponding to the breathing chimera (red lines in Fig. 2(b), yellow-shaded region). Fig. 3(a) shows exemplary time series of this state, and Fig.3(c) the corresponding phase portrait, with black curves depicting the limit-cycles and the red curve showing an example of a trajectory.
@@
    The breathing chimera limit-cycles eventually collide with the unstable coherent homogeneous state at $K\approx 7.89$, merging in a single limit-cycle through a double homoclinic bifurcation (Dhom., also known as gluing bifurcation Lopez and Marques (2000); Pazó and Pérez-Muñuzuri (2001); Laing (2012)). The resulting attractor corresponds to a symmetric limit cycle in which the two populations alternate the same level of synchrony half a period apart (red lines in Fig. 2(b), green-shaded region). Figure 3(b) shows exemplary time series of this regime and Figs. 3(d) and (e) depict two phase portraits, with $K$ in panel (d) very close to the gluing bifurcation.
@@
    At $K\approx 15.67$ the homogeneous fixed point recovers stability through the (subcritical) pitchfork bifurcation provided by Eq. (14) (see PF- in Fig. 2(c)). Therefore, in this region both homogeneous and antiphase states share stability (black curves in Fig. 2(c)).
@@
    Finally, the unstable fixed points emerging from PF- collide with the antiphase solution at a new subcritical pitchfork bifurcation (PF’-) at $K\approx 33.69$ (see Appendix D for an analytical derivation of PF’-). From this point, further increase of $K$ does not lead to any other bifurcation, and the homogeneous state remains the only (global) attractor of the system (black curve in Fig. 2(c)).
@@
The two-parameter bifurcation diagram in Fig. 2(a) shows that the bifurcation scenario we have described for $\mu=0.9$ holds for the entire region of instability of the homogeneous state. The diagram also shows the loci of a saddle node bifurcation (SN) that joints the pitchfork bifurcation (PF) at a fold-pitchfork codimension-2 bifurcation (SN-PF). At this codimension 2 point, the symmetry breaking bifurcation turns from supercritical to subcriticalLaing (2009a); Kotwal, Jiang, and Abrams (2017). This SN branch limits a region of bistability between the homogeneous fixed point and stable chimeras. Within this region, following a route similar to the one described before, the stable chimera turns to a breathing state at the supercritical Hopf (H+). Nonetheless, in this case the gluing bifurcation leads to the disappearance of limit-cycle solutions, since the homogeneous state is stable. Finally, there is an additional region of bistability between the asymmetric limit cycle and the homogeneous state bounded by a heteroclinic bifurcation and the saddle-node of limit cycles (SNLC).
Altogether, this analysis unveils a rich dynamical landscape emerging from the pitchfork bifurcation breaking the symmetry of the homogeneous state. These results complement those presented in previous works for the same system Laing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021) but without relying on weak heterogeneity. Morevoer, the pitchfork bifurcation of the homogeneous state (PF) and the bifurcations from the antiphase state (H- and PF’-) have been obtain analytically as an implicit equation on the system parameters.
@@ V Ring topology
This scenario was studied in Ref. [Lee and Krischer, 2023b] for the case of $N=6$ oscillators, where they unveiled different forms of stable and (non-chaotic) breathing chimera for heterogeneous populations. Also, Ref. [Laing, 2023] analyzed a similar ring model but with non-local coupling, which displayed chaotic solutions that vanish upon increasing system size. Here we show that, in our setup, transverse instabilities of the homogeneous state lead to chaotic dynamics that show an extensive character with $N$, i.e., the dynamical complexity of the system scales linearly with the system size.
@@
As $N$ increases, the second largest structural eigenvalue $\Lambda_{2}$ tends to 1. Therefore, from Fig. 1(b), we infer that the region of transverse instability expands with $N$. Fig. 4(a) illustrates this situation: Thick continuous curves depict the bifurcation obtained from setting $\text{Re}[\lambda^{(+)}]=0$ in Eq. (14) in the $(K,\mu)$-plane. As $N$ increases, the region of instability becomes larger, covering a wide band of $K$ values irrespective of $\mu$ if $N$ is sufficiently large. Moreover, the emergence of instabilities occurs arbitrarily close to the Kuramoto synchronization transition (H0, vertical black dashed line in Fig. 4(a)).
Another important aspect of the ring spectrum in Eq. (18) is that it is delocalized, i.e., it covers the interval $\Lambda_{k}\in[2\mu-1,1]$ densely as $N\to\infty$. Therefore, several structural eigenmodes $\Lambda_{k}$ might correspond to unstable directions concurrently. For instance, the colored region in Fig. 4(b) shows the number of unstable directions of the homogeneous state in the $(K,\mu)$-plane for $N=128$. The number of positive growth rates generally increases with $\mu$, with a large region (dark blue) where all $\lambda_{k}^{(+)}>0$ for $k>1$. In fact, this region coincides with the region of transverse instability for the two-population model, since the last mode turning unstable corresponds to $\Lambda_{N}=2\mu-1$.
@@
The regularity of ring topologies also allows for the existence of antiphase and other twisted states Lee and Krischer (2023b). Appendix D provides a linear stability analysis for antiphase states in ring networks. Thin lines in Fig. 4(a) and red curve in (b) delimit the region of stable antiphase dynamics. This region shrinks by increasing $N$, although the changes are less prominent than those of the homogeneous state (see Appendix D for a further discussion).
Overall, in large ring networks, we expect to have complex dynamics in wide regions of the parameter space due to the expansion of the transverse instability region and the number of unstable directions of the homogeneous state. Figure 5 shows two instances of these complex dynamics for $N=128$. Pannels (a) depict $R_{\sigma}$ and $\Phi_{\sigma}$ of each population over time for parameters in the region where the number of unstable directions of the homogeneous state is maximal ($K=7$, $\mu=0.9$). The dynamics of $R_{\sigma}$ are rather hectic, with $\phi_{\sigma}$ showing some indications of spatial structure. Instead, panels (b) correspond to parameters associated with less unstable directions ($K=15$, $\mu=0.5$). In this case, the dynamics of both $R_{\sigma}$ and $\Phi_{\sigma}$ evolve more regularity in spite of their seemingly chaotic behavior. These qualitative observations posit the existence of space-time chaos emerging from the transverse instabilities. In the rest of this section, we asses the chaotic and complex behavior of the system by means of numerical Lyapunov analysis. Appendix A contains precise definitions of the monitored quantities.
First, we focus on a fixed network size ($N=128$) and study the dynamics by varying $K$. Figures 6(a1-d1) show the numerical results obtained with fixed $\mu=0.9$, thus covering the region of maximal number of unstabilities. Red circles in Figure 6(a1) show the time average of the network’s mean Kuramoto order parameter, $\mathcal{R}$. The shaded region shows the corresponding standard deviation $\mathcal{S}$ (see Appendix A for specific definitions). Grey continuous lines indicate the value of $\mathcal{R}$ corresponding to the homogeneous and antiphase states (Eqs. (5) and (29)). As predicted, the nontrivial dynamics emerge just after the Kuramoto transition to synchrony $H_{0}$ and appear to be the only attracting state until the stabilization of the antiphase state (see dotted vertical lines).
@@
Next, we switch to $\mu=0.5$, for which the homogeneous state has less unstable directions. Moreover, the antiphase state does not exist for this value of $\mu$, and therefore, we expect non-trivial dynamics for a larger $K$-range. Figures 6(a2-d2) show that the scenario starts off analogously to that of $\mu=0.9$: at the transverse instability (see vertical dotted line) the system transitions from incoherence to a state with a large degree of heterogeneity. Also, several Lyapunov exponents become positive at once, resulting into both $\mathcal{D}_{\text{KY}}$ and $h_{\text{KS}}$ increasing rapidly. Again, this corresponds to a case of space-time chaos. Nonetheless, around $K\approx 15.8$ a transition occurs in which the irregularity of the system drops drastically.
The dynamics emerging from this bifurcation, which is not detected by our theory, evolve very close to the homogeneous state. In fact, our measure of network heterogeneity $\mathcal{S}$ does not capture spatial inhomogeneities (see red symbols in Fig.6(a2)). Nonetheless, a close inspection (not shown) revealed irregular space-time patterns in both, $R_{\sigma}$ and $\Phi_{\sigma}$, but the dynamics remains very close to the homogeneous solution and evolve rather slowly compared to the previous chaotic regime. The most clear indication of irregularity are the larger Lyapunov exponents, which are closer to zero, but many of them are still positive according to our numerics (see panels (b2-d2) in Fig.6). Finally, as shown by $\mathcal{D}_{\text{KY}}$, the dynamics fall back to the homogeneous state slightly after the predicted transition ($K\approx 21$, vertical dashed line in 6). This small offset between our analysis and the numerics is due to the slow convergence of small negative Lyapunov exponents in this regime.
@@
So far, the numerical exploration for $\mu=0.9$ and $\mu=0.5$ proved the existence of highly complex spatiotemporal regimes emerging from transverse instabilities of the homogeneous state in rings of $N=128$ populations. Next, we test the robustness and scaling these results upon varying the system size $N$. Fig. 7 shows results from simulations of rings composed of $N=64,128,256,512$ and $1024$ populations (red and blue symbols) for selected parameter values. Panel (a) shows the Lyapunov spectra computed for different system sizes with normalized node index. For the tested parameter values ($(K,\mu)=(7,0.9)$ and $(15,0.5)$) the spectra shows a good collapse upon increasing $N$, with the only exception of the smaller networks for the second set (blue dots). This robust spectra upon increasing $N$ indicates extensive chaos, i.e., a chaotic regime with complexity scaling linearly with the system size. To further explore this possibility, Figs. 7(b) and (c) show the computed fractal dimension and dynamical entropy for the same parameter values upon increasing $N$. In all cases, an accurate scaling $\mathcal{D}_{\text{KY}}\propto N$ and $h_{\text{KS}}\propto$ is observed.
@@
Figure 6: Simulations of ring (columns 1 and 2) and Erdős-Rényi (column 3) networks with $N=128$ ($\alpha=1.2$ and $\Delta=1$). (a) Red circles indicate time average mean activity $\langle\mathcal{R}\rangle$, with the red-shaded region showing the corresponding level of heterogeneity ±$\langle\mathcal{R}\rangle\pm\langle\mathcal{S}\rangle$. Grey curves correspond to the value of $R$ for the homogeneous states given by Eq. (5). In panel (a1) an additional grey curve shows the antiphase state (Eq. (29)). (b) 20 largest Lyapunov exponents $\ell_{\sigma}$. (c) Kaplan-Yorke fractal dimension computed using the (complete) Lyapunov spectra as given by Eq. (23). (d) Kolmogorov-Sinai dynamical entropy computed from Eq. (24). In all panels, vertical dashed lines indicate the emergence of transverse instabilities as given by Eq. (14). Dotted vertical line in panels (a1-d1) indicates the emergence of stable antiphase state as given by Eq. (35).
Figure 7: Extensivity of chaos in ring and Erdős-Rényi (ER) networks ($\alpha=1.2$ and $\Delta=1$). (a) Complete Lyapunov spectra for different network sizes for the simulations of the ring and ER networks. For each parameter set, dots indicate the results for networks with $N=64,128,256,512$ and 1024 populations. Nonetheless, in most cases differences across system sizes are barely visible. (b,c) Fractal dimension $D_{\textbf{KY}}$ (panel b) and dynamical entropy (panel c) $h_{\textbf{KS}}$ computed for different system sizes $N$ in three different simulation setups. Circles correspond to numerical simulations. Dashed lines correspond to a linear regression. In all these plots, red corresponds to ring networks with $K=7$ and $\mu=0.9$, blue corresponds to ring networks with $K=15$ and $\mu=0.5$, and green corresponds to ER networks with $K=14$ and $\mu=0.9$.

### v2

@@ IV Review of the two-population model
Figure 2: Bifurcation diagrams of the two population model (Eq. (3) with Eq. (16)) for $\alpha=1.2$ and $\Delta=1$. (a) Two-parameter bifurcation diagram. Solid lines indicate supercritical bifurcations (+) and dashed lines indicate subcritical bifurcations (-). Black curves correspond to the pitchfork bifurcation (PF) of the homogeneous state (given by Eq. (14)), red curves indicate Hopf bifurcations (H), green curves indicate double homoclinic (DHom.) and heteroclinic (Het.) bifurcations, the blue curve indicates a saddle-node bifurcation (SN), and the purple curve indicates a saddle-node of limit-cycles (SNLC). Straight dotted thin line indicates the Kuramoto synchronization transition (H0). Color-shaded regions indicate instability of the homogeneous solution, with each color corresponding to a different attractor set: pink for stable steady chimeras, yellow for breathing chimeras, green for symmetric oscillations, blue for bistability between symmetric oscillations and antiphase state, and purple for stability of the antiphase fixed point only. (b,c) One-parameter bifurcation diagram for $\mu=0.9$. Solid curves with different colors indicate different attractor types: black for stable fixed points, grey for unstable fixed points, red for stable limit-cycles, and light red for unstable limit-cycles. Relevant bifurcations are labelled on the top of the figures, with the different regions between bifurcations shaded with the same color scheme as in panel (a). Homogeneous and antiphase solutions together with their corresponding bifurcations (H0,PF,H-) have been obtained analytically. All other results have been obtained with auto-07pDoedel et al. (2007) (see Appendix B).
As a first example we revisit the two-population model. Previous works on this setup showed the emergence of chimera and other states from a pitchfork bifurcation of the homogeneous synchronized state if $\Delta>0$Laing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021). Here we show that such bifurcation corresponds to a transverse instability provided by Eq. (14), and review some of the dynamics emerging from this symmetry-breaking. As in the previous section, we fix $\Delta=1$ without loss of generality (see Appendix C).
@@
The eigenvalues of this connectivity matrix read $\Lambda_{1}=1$ and $\Lambda_{2}=2\mu-1$. Substituting $\Lambda_{2}$ in Eq. (14) and setting $\text{Re}[\lambda_{2}^{(+)}]=0$ provides an analytical expression for the pitchfork bifurcation of the homogeneous state as an implicit equation on the system parameters. The solid black curve in Figure 2(a) shows such bifurcation line in the $(K,\mu)$-plane, with the color-shaded regions above the line indicating the areas where the homogeneous state is unstable (see also black curve in Fig. 1(b)). Within this region, four different states might exist (apart from the incoherent and coherent homogeneous states) [^5] For $\alpha=1.2$ we did not find the chaotic state unveiled in Ref. [Laing, 2012], but it can be attained, for instance, at $(K,\mu,\alpha,\Delta)=(20,0.9,1.25,1)$:
  • Steady chimera (pink region in Fig. 2a), a fixed point characterized by $R_{1}\neq R_{2}$.
  • Breathing chimera (yellow region in Fig. 2a), a limit-cycle characterized by one population having always a larger degree of synchrony than the other, i.e., either $R_{1}(t)>R_{2}(t)$ or $R_{2}(t)>R_{1}(t)$ for all $t>0$ (see Fig. 3(a)).
  • Symmetric limit-cycle (green and blue regions in Fig. 2a), characterized by $R_{1}(t)=R_{2}(t+T/2)$, where $T$ is the period of the oscillation (see Fig. 3(b)).
  • Antiphase fixed point (blue and purple regions in Fig. 2a), characterized by $R_{1}(t)=R_{2}(t)$ but $\Phi_{1}-\Phi_{2}=\pi$. A thorough analysis of this state, including its linear stability, is provided in Appendix D.
In order to illustrate the emergence and nature of these different states, we first fix $\mu=0.9$ and study the system dynamics upon increasing $K$ with the help of numerical continuation softwareDoedel et al. (2007). The resulting bifurcation diagram is depicted in Figs. 2(b,c), and can be split in nine different regions (labelled i-ix) delimited by different bifurcations:
@@
    For $K<K_{c}\approx 5.52$ the incoherent state $Z_{\sigma}=0$ is the only stable solution. This state loses stability at the Kuramoto synchronization transition H0, leading to the emergence of the coherent homogeneous state.
@@
    The coherent homogeneous state is the global attractor of the system from H0 up to $K\approx 6.66$, when it loses stability through the supercritical pitchfork bifurcation (PF+) given by Eq. (14).
@@
    Above PF+, two stable fixed points emerge, corresponding to steady chimeras (black lines in Fig. 2(b), red-shaded region). Although the steady chimeras coexist with the unstable homogeneous and antiphase fixed points, they are the sole attractor in this region of the parameter space.
@@
    At $K\approx 7.37$ the steady chimeras undergo a supercritical Hopf bifurcation (H+) leading to the emergence of two stable limit-cycles, corresponding to the breathing chimera (red lines in Fig. 2(b), yellow-shaded region). Figure 3(a) shows exemplary time series of this state, and Fig. 3(c) the corresponding phase portrait, with black curves depicting the limit-cycles and the red curve showing an example of a trajectory.
@@
    The breathing chimera limit-cycles eventually collide with the unstable coherent homogeneous state at $K\approx 7.89$, merging in a single limit-cycle through a double homoclinic bifurcation (Dhom, also known as gluing bifurcation Lopez and Marques (2000); Pazó and Pérez-Muñuzuri (2001); Laing (2012)). The resulting attractor corresponds to a symmetric limit cycle in which the two populations alternate the same level of synchrony half a period apart (red lines in Fig 2(b), green-shaded region). Figure 3(b) shows exemplary time series of this regime and Figs. 3(d,e) depict two phase portraits, with $K$ in panel (d) very close to the gluing bifurcation.
@@
    At $K\approx 15.67$, the homogeneous fixed point recovers stability through the (subcritical) pitchfork bifurcation provided by Eq. (14) (see PF- in Fig. 2(c)). Therefore, in this region both homogeneous and antiphase states share stability (black curves in Fig. 2(c)).
@@
    Finally, the unstable fixed points emerging from PF- collide with the antiphase solution at a new subcritical pitchfork bifurcation (PF’-) at $K\approx 33.69$ (see Appendix D for an analytical derivation of PF’-). From this point on, further increase of $K$ does not lead to any other bifurcation, and the homogeneous state remains the only (global) attractor of the system (black curve in Fig. 2(c)).
@@
The two-parameter bifurcation diagram in Fig. 2(a) shows that the bifurcation scenario we have described for $\mu=0.9$ holds for the entire region of instability of the homogeneous state. The diagram also shows the loci of a saddle node bifurcation (SN) that joints the pitchfork bifurcation (PF) at a fold-pitchfork codimension-2 bifurcation (SN-PF). At this codimension 2 point, the symmetry breaking bifurcation turns from supercritical to subcriticalLaing (2009a); Kotwal, Jiang, and Abrams (2017). This SN branch limits a region of bistability between the homogeneous fixed point and stable chimeras. Within this region, following a route similar to the one described before, the stable chimera turns into a breathing state at the supercritical Hopf (H+). Nonetheless, in this case the gluing bifurcation leads to the disappearance of limit-cycle solutions, since the homogeneous state is stable. Finally, there is an additional region of bistability between the asymmetric limit cycle and the homogeneous state bounded by a heteroclinic bifurcation and the saddle-node of limit cycles (SNLC).
Altogether, this analysis unveils a rich dynamical landscape emerging from the pitchfork bifurcation that breaks the symmetry of the homogeneous state. These results complement those presented in previous works for the same system Laing (2009a, 2012); Kotwal, Jiang, and Abrams (2017); Guo et al. (2021) but without relying on weak heterogeneity. Moreover, the pitchfork bifurcation of the homogeneous state (PF) and the bifurcations from the antiphase state (H- and PF’-) have been obtained analytically as implicit equations on the system parameters.
@@ V Ring topology
This scenario was studied in Ref. [Lee and Krischer, 2023b] for the case of $N=6$ oscillators, where they unveiled different forms of stable and (non-chaotic) breathing chimera states for heterogeneous populations. Also, Ref. [Laing, 2023] analyzed a similar ring model but with non-local coupling, which displayed chaotic solutions that vanish upon increasing system size. Here we show that, in our setup, transverse instabilities of the homogeneous state lead to chaotic dynamics that show an extensive character with $N$, i.e., the dynamical complexity of the system scales linearly with the system size.
@@
As $N$ increases, the second largest structural eigenvalue $\Lambda_{2}$ tends to 1. Therefore, from Fig. 1(b), we infer that the region of transverse instability expands with $N$. Figure 4(a) illustrates this situation: Thick continuous curves depict the bifurcation obtained in the $(K,\mu)$-plane by setting $\text{Re}[\lambda^{(+)}]=0$ in Eq. (14). As $N$ increases, the region of instability becomes larger, covering a wide band of $K$ values irrespective of $\mu$ if $N$ is sufficiently large. Moreover, the emergence of instabilities occurs arbitrarily close to the Kuramoto synchronization transition (H0, vertical black dashed line in Fig. 4(a)).
Another important aspect of the ring spectrum in Eq. (18) is that it is delocalized, i.e., it covers the interval $\Lambda_{k}\in[2\mu-1,1]$ densely as $N\to\infty$. Therefore, several structural eigenmodes $\Lambda_{k}$ might concurrently correspond to different unstable directions. For instance, the colored region in Fig. 4(b) shows the number of unstable directions of the homogeneous state in the $(K,\mu)$-plane for $N=128$. The number of positive growth rates generally increases with $\mu$, with a large region (dark blue) where all $\lambda_{k}^{(+)}>0$ for $k>1$. In fact, this region coincides with the region of transverse instability for the two-population model, since the last mode turning unstable corresponds to $\Lambda_{N}=2\mu-1$.
@@
The regularity of ring topologies also allows for the existence of antiphase and other twisted states Lee and Krischer (2023b). Appendix D provides a linear stability analysis for antiphase states in ring networks. The thin lines in Fig. 4(a) and the red line in Fig. 4(b) delimit the region of stable antiphase dynamics. This region shrinks by increasing $N$, although the changes are less prominent than those of the homogeneous state (see Appendix D for a further discussion).
Overall, in large ring networks, we expect to have complex dynamics in wide regions of parameter space due to the expansion of the transverse instability region and the number of unstable directions of the homogeneous state. Figure 5 shows two instances of these complex dynamics for $N=128$. Pannels (a) depict $R_{\sigma}$ and $\Phi_{\sigma}$ of each population over time for parameters in the region where the number of unstable directions of the homogeneous state is maximal ($K=7$, $\mu=0.9$). The dynamics of $R_{\sigma}$ are rather disordered, with $\phi_{\sigma}$ showing some indications of spatial structure. Instead, panels (b) correspond to parameters associated with less unstable directions ($K=15$, $\mu=0.5$). In this case, the dynamics of both $R_{\sigma}$ and $\Phi_{\sigma}$ evolve more regularly, in spite of their seemingly chaotic behavior. These qualitative observations posit the existence of space-time chaos emerging from the transverse instabilities reported here. In the rest of this section, we assess the chaotic and complex behavior of the system by means of numerical Lyapunov analysis. Appendix A contains precise definitions of the monitored quantities.
First, we focus on a fixed network size ($N=128$) and study the dynamics by varying $K$. Figures 6(a1-d1) show the numerical results obtained with fixed $\mu=0.9$, thus covering the region of maximal number of unstabilities. The red circles in Fig. 6(a1) show the time average of the network’s mean Kuramoto order parameter, $\mathcal{R}$. The shaded region shows the corresponding standard deviation $\mathcal{S}$ (see Appendix A for specific definitions). The grey continuous lines indicate the value of $\mathcal{R}$ corresponding to the homogeneous and antiphase states (Eqs. (5) and (29)). As predicted, the nontrivial dynamics emerge just after the Kuramoto transition to synchrony H0, and appear to be the only attracting state until the stabilization of the antiphase state (see dotted vertical lines).
@@
Next, we switch to $\mu=0.5$, for which the homogeneous state has less unstable directions. Moreover, the antiphase state does not exist for this value of $\mu$, and therefore, we expect non-trivial dynamics for a larger $K$-range (note the difference in the x-axis range between columns 1 and 2 in Fig. 6). Figures 6(a2-d2) show that the scenario starts off analogously to that of $\mu=0.9$: at the transverse instability (see vertical dotted line) the system transitions from incoherence to a state with a large degree of heterogeneity. Also, several Lyapunov exponents become positive at once, resulting into both $\mathcal{D}_{\text{KY}}$ and $h_{\text{KS}}$ increasing rapidly. Again, this corresponds to a case of space-time chaos.
Around $K\approx 15.8$ a transition occurs and the irregularity of the system drops drastically. The dynamics emerging from this bifurcation, which is not detected by our theory, proved challenging to characterize. First, the system suddenly evolves very close to the homogeneous state. In fact, our measure of network heterogeneity $\mathcal{S}$ is too small to be visually appreciated above the transition in Fig. 6(a2) (see pink shaded region). Nonetheless, a close inspection of the simulations (not shown) revealed irregular patterns in both $R_{\sigma}$ and $\Phi_{\sigma}$. Second, such spatiotemporal dynamics evolve slowly compared to the states encountered so far, requiring longer simulations (see Appendix B). Such slow time-scales are mainly reflected by the magnitude of the positive Lyapunov exponents in Fig. 6(b2) and the drop on dynamical entropy shown in Fig. 6(d2). In spite of these peculiarities, the number of positive Lyapunov exponents, and thus the value of $\mathcal{D}_{\text{KY}}$ for $15.8\lesssim K\lesssim 21$ (see Fig. 6(c3)) indicate that this corresponds to a high-dimensional chaotic regime  [^6] We also performed simulations with adiabatic increase and decrease of $K$ in other to test for potential bistability between the fast and slow chaotic states. The outcome consistently showed the disappearance of the faster dynamics around $K\approx$.. Finally, around $K\approx 21.6$ the dynamics fall back to the homogeneous state, as predicted by the stability analysis (second vertical dashed line in the middle-row panels of Fig. 6).
@@
So far, the numerical exploration for $\mu=0.9$ and $\mu=0.5$ proved the existence of highly complex spatiotemporal regimes emerging from transverse instabilities of the homogeneous state in rings of $N=128$ populations. Next, we test the robustness and scaling these results upon varying the system size $N$. Figure 7 shows results from simulations of rings composed of increased number $N$ of oscillators (red and blue symbols) for selected parameter values. Panel (a) shows the Lyapunov spectra computed for different system sizes with node index normalized by the system size $\sigma/(2N)$. For the tested parameter values ($(K,\mu)=(7,0.9)$, red symbols, and $(15,0.5)$, blue symbols) the spectra show a good collapse upon increasing $N$, i.e., the Lyapunov exponents converge to a well defined profile. The only exception of the smaller networks for the second set (blue dots). These robust spectra for increasing $N$ indicate extensive chaos, i.e. a chaotic regime whose complexity scales linearly with the system size. To further explore this possibility, Figs. 7(b) and (c) show the computed fractal dimension and dynamical entropy for the same parameter values upon increasing $N$. In all cases, an accurate scaling $\mathcal{D}_{\text{KY}}\propto N$ and $h_{\text{KS}}\propto N$ is observed.
@@
Figure 6: Simulations of ring (columns 1 and 2) and Erdős-Rényi (column 3) networks with $N=128$ ($\alpha=1.2$ and $\Delta=1$). (a) Red circles indicate time average mean activity $\langle\mathcal{R}\rangle$, with the red-shaded region showing the corresponding level of heterogeneity ±$\langle\mathcal{R}\rangle\pm\langle\mathcal{S}\rangle$. Grey curves correspond to the value of $R$ for the homogeneous states given by Eq. (5). In panel (a1) an additional grey curve shows the antiphase state (Eq. (29)). (b) 20 largest Lyapunov exponents $\ell_{\sigma}$. (c) Kaplan-Yorke fractal dimension computed using the (complete) Lyapunov spectra as given by Eq. (23). (d) Kolmogorov-Sinai dynamical entropy computed from Eq. (24). In all panels, vertical dashed lines indicate the emergence and disappearance of transverse instabilities as given by Eq. (14). Dotted vertical line in panels (a1-d1) indicates the emergence of stable antiphase state as given by Eq. (35).
Figure 7: Extensivity of chaos in ring and Erdős-Rényi (ER) networks ($\alpha=1.2$ and $\Delta=1$). (a) Complete Lyapunov spectra for different network sizes for the simulations of the ring and ER networks. For each parameter set, dots indicate the results for networks with $N=64,128,256,512$ and 1024 populations. Nonetheless, in most cases differences across system sizes are barely visible. (b,c) Fractal dimension $D_{\text{KY}}$ (panel b) and dynamical entropy (panel c) $h_{\text{KS}}$ computed for different system sizes $N$ in three different simulation setups. Circles correspond to numerical simulations. Dashed lines correspond to a linear regression. In all these plots, red corresponds to ring networks with $K=7$ and $\mu=0.9$, blue corresponds to ring networks with $K=15$ and $\mu=0.5$, and green corresponds to ER networks with $K=14$ and $\mu=0.9$.